# Reducing Numerical Precision Requirements in Quantum Chemistry Calculations

### v1

@@ IV.2 Ozaki Scheme Application
Despite this promising finding, it is unlikely that future hardware will offer implementations of floating point types with effective 37 bit mantissas. Instead, we propose using the Ozaki scheme with a reduced number of splits to take advantage of the reduced precision needs. In Fig. 7, we plot the errors as a function of the number of splits for various floating point representations. When using NVIDIA’s Tensor Cores (FP16 with FP32 accumulation), four to five splits (10 to 15 multiplications) is sufficient for a converged result; the difference the absolute error in the energy between 4 and 5 splits is $2.94\times 10^{-10}$ Hartree ($1.34\times 10^{-8}$ Hartree for three and four splits). By contrast, according to Fig. 3 to fully reproduce a double precision result on random matrices required 6 splits / 21 multiplications. Hence, the Ozaki scheme allows for approximately a factor of two savings in computational cost due to the lower precision requirements for quantum chemistry application.
@@
Figure 7: Error in the density matrix when using purification vs. the number of splits used in the Ozaki scheme applied to the Molnupiravir / BigDFT system (matrix dimension $147\times 147$). We examine the error for four different floating point types: BFLOAT16 (8, 7), FP16 (5, 10), AMD FP24 (7, 16), and FP16 with FP32 (8, 23) Accumulation (TC).
@@
In all cases, the scaling scheme seamlessly handles the exponent, and hence the precision comes down to the size of the mantissa. Using plain FP16, 16 - 18 splits (136 - 171 multiplications) is required. This is a substantial increase in the number of multiplications compared to the Tensor Core version, however future hardware may not offer FP16 TC (especially non-NVIDIA chips), with different performance ratios. For a fully converged double precision result on random matrices, 20 splits / 210 multiplications was required, leading again to an approximate factor of two savings. We compare as well the intermediate sized mantissa of AMD’s FP24, and find that it converges around 7 splits / 28 multiplications. For this system, the TRS4 algorithm required 26 calls to the Ozaki scheme multiplication routine to converge. Hence, when using Tensor Cores an overall speedup would require a time to solution ratio of 260 – 390 multiplications to one double precision diagonalization. For pure FP16, the requirement increases to 3536 multiplications.
IV.3 Potential to Exploit Sparsity
@@
Depending on the underlying distribution of matrix values, the splitting procedure may introduce a significant amount of sparsity. This can potentially be used to accelerate the calculation: on an A100 GPU, the FP16 tensor core performance increases from 312 TFLOPS to 624 TFLOPS for sparse matrices. We examined the final multiplication performed for the Molnupiravir system by the purification algorithm when using the Tensor Core representation (FP16 with FP32 accumulation) and 5 splits. An entry of the split matrix (after scaling) was considered to be zero if it fell below the smallest value representable by double precision. The first split of both matrices have a number of non-zeros below 40%, however the next split is above 94%, and subsequent splits are nearly fully dense. For the pure FP16 implementation with 16 splits, the matrices are substantially more sparse: the number of non-zeros is below 3% for the first split, 7% for the second, and 14% for the third split. By the sixth split the sparsity of both matrices was above 50% (and for all subsequent splits). Thus we anticipate that some sparsity may be exploited in future work, though it would be limited to the first set of splits.
IV.4 Basis Set and Size Effects
@@
Figure 8: Impact of the basis set on the number of splits required in the Ozaki scheme for the selenite in water system. Matrix sizes are: $179\times 179$ (PCSEG-0), $309\times 309$ (PCSEG-1), $515\times 515$ (AUG-PCSEG-1), $713\times 713$ (PCSEG-2), $1117\times 1117$ (AUG-PCSEG-2), and $1508\times 1508$ (PCSEG-3). All calculations are performed with FP16 Tensor Cores.
In Fig. 8 we plot the error for each of the basis sets with a given number of splits. We observe that for the larger basis sets, the required number of splits increases from four to five; with five splits, the error in the norm for PCSEG-3 is $1.1\times 10^{-6}$ and absolute error in energy is $3.2\times 10^{-8}$ Hartree. This is consistent with our previous results where the difference between four and five splits yield was small, but not non-negligible. Overall, we find that our earlier findings transfer well to even challenging numerical conditions. Thus, purification with the Ozaki scheme should be applicable to routinely performed quantum chemistry calculations, and not just specialized approximate schemes.
Figure 9: Impact of the system size on the number of splits required in the Ozaki scheme for the water cluster systems. BigDFT uses 6 basis functions per water molecule, so that the smallest matrix is of dimension $30\times 30$ and the largest $1806\times 1806$. All calculations are performed with FP16 Tensor Cores.
We also investigate the number of splits required as a function of system size using water clusters of increasing size computed with BigDFT. This analysis will allow us to separate the effect of basis set conditioning with matrix size. For the water clusters, the largest matrix is of dimension $1806\times 1806$, and for the selenite systems it was $1508\times 1508$ (PCSEG-3). The errors with respect to system size are plotted in Fig. 9. We find a modest growth in the number of splits required with system size. The error with five splits is similar for the largest water cluster as for the selenite system with the largest basis set; the error in the norm $5.2\times 10^{-7}$ is and absolute error in energy is $5.8\times 10^{-8}$ Hartree. Thus, we conclude that five splits is a sufficient recommendation, though automated schemes may improve the usability of the Ozaki scheme.
IV.5 Comparison with the Markidis Method
In the previous work of Finkelstein et al. that performed density matrix purification on NVIDIA Tensor Core units [39, 40, 41], the precision of the result was improved using the method of Markidis [2]. Here, the matrix $X$ is split into a lower and high precision part:
@@ IV.5 Comparison with the Markidis Method
In the Markidis method, we can include up to four terms, each refining the precision of the result. In Tab. 1 we show the precision and the number of purification iterations required (which may increase due to low precision) for a given number of terms for various systems. We caution that the number of iterations may be made more uniform by a more sophisticated convergence test [58].
@@
Errors
Molnupiravir $1.6\times 10^{-2}$ $1.0\times 10^{-2}$ $2.1\times 10^{-5}$ $3.1\times 10^{-5}$ $8.4\times 10^{-6}$ $6.3\times 10^{-11}$ $5.1\times 10^{-13}$
Water 237 $3.9\times 10^{-2}$ $2.5\times 10^{-2}$ $1.4\times 10^{-4}$ $4.3\times 10^{-4}$ $3.2\times 10^{-5}$ $3.0\times 10^{-7}$ $5.7\times 10^{-12}$
Selenite/A $8.7\times 10^{-2}$ $4.9\times 10^{-2}$ $1.7\times 10^{-4}$ $1.6\times 10^{-4}$ $4.9\times 10^{-5}$ $1.0\times 10^{-8}$ $1.7\times 10^{-13}$
Selenite/B $9.7\times 10^{-2}$ $5.8\times 10^{-2}$ $3.1\times 10^{-4}$ $3.3\times 10^{-4}$ $1.6\times 10^{-4}$ $2.0\times 10^{-8}$ $3.4\times 10^{-13}$
Selenite/C $2.1\times 10^{-1}$ $1.2\times 10^{-1}$ $2.1\times 10^{-3}$ $2.2\times 10^{-3}$ $4.9\times 10^{-4}$ $1.1\times 10^{-6}$ $7.8\times 10^{-13}$
@@
Molnupiravir $17$ $27$ $16$ $23$ $16$ $13$ $13$
Water 237 $10$ $13$ $17$ $15$ $15$ $11$ $11$
Selenite/A $100$ $32$ $22$ $29$ $22$ $21$ $21$
Selenite/B $63$ $100$ $37$ $33$ $27$ $27$ $27$
Selenite/C $100$ $66$ $31$ $41$ $44$ $31$ $28$
@@
Table 1: Errors in the Frobenius norm of the density matrix and number of purification iterations for the various systems computed with FP16 Tensor Cores vs. the numbers of correction terms. A: PCSEG-1; B: AUG-PCSEG-1; C: PCSEG-3. 1 Term: $A^{0}B^{0}$; 2 Terms: $A^{0}B^{0}+A^{0}B^{1}$; 3 Terms: $A^{0}B^{0}+A^{0}B^{1}+A^{1}B^{0}$; 4 Terms: $A^{0}B^{0}+A^{0}B^{1}+A^{1}B^{0}+A^{0}B^{0}$. 4-MPFR: the 4 Term MPFR implementation with infinite exponent. Ozaki (5): Ozaki scheme with 5 splits. FP64: calculations performed with double precision.
@@
While comparing our CUDA and MPFR implementations of the Markidis method, we noted that a significant amount of error was introduced due to the limited exponent range. To improve this, we applied the same scaling method for the exponent ranges used in the Ozaki scheme [44], and found it provided substantial improvement. For example, the 4 Terms result for Selenite/PCSEG-3 improved from an error of $1.30\times 10^{-2}$ to $2.16\times 10^{-3}$. Nonetheless, the MPFR result still remains more precise; in the future, it may be possible to further improve the Markidis method with a new scaling scheme.
@@
In the implementation of Finkelstein et al., they include three terms in the Markidis correction, which is well justified from our data. This three term result can be accomplished at the cost of only two multiplications because they implement a scheme that only requires squaring symmetric matrices, which means that one can exploit the relation $A^{1}A^{0}=(A^{0}A^{1})^{T}$ even in inexact arithmetic. The benefit of the Markidis method is thus that it can substantially improve the accuracy at a low cost with FP16 Tensor Cores (especially if combined with the exponent scaling scheme). On the other hand, it does not converge to the double precision result, can’t take advantage of lower precision hardware like FP16 without FP32 accumulation, and has significant errors in more challenging numerical conditions (like larger systems and basis sets). For example, for the 3 Term scheme the absolute error in energy for Molnupiravir is $4.4\times 10^{-5}$ Hartree, but it is $6.5\times 10^{-3}$ Hartree for selenite with the PCSEG-3 basis set. For this reason, the higher order approximation of the Ozaki scheme is valuable. When only the matrix square is required, the Ozaki scheme can also exploit this symmetry in the calculation of the error free terms (see Fig. 2) as well as other optimizations detailed by Uchino [59].
IV.6 Mixed Precision Purification

### v2

@@ 4.2 Ozaki Scheme Application
Despite this promising finding, it is unlikely that future hardware will offer implementations of floating point types with effective 37 bit mantissas. Instead, we propose using the Ozaki scheme with a reduced number of splits to take advantage of the reduced precision needs. In Fig. 7, we plot the errors as a function of the number of splits for various floating point representations. When using NVIDIA’s Tensor Cores (FP16 with FP32 accumulation), three to four splits (6 – 10 multiplications) is sufficient for a converged result. By contrast, according to Fig. 3 to fully reproduce a double precision result on random matrices required 6 splits / 21 multiplications. Hence, the Ozaki scheme allows for approximately a factor of two savings in computational cost due to the lower precision requirements for quantum chemistry application. These multiplication counts can be compared with Tab. 1, where we see that FP16 TCs can have above 15 times the performance of FP64 TCs, and over a factor of 100 more performance for standard double precision operations.
@@
Figure 7: Errors when using purification vs. the number of splits used in the Ozaki scheme applied to the Molnupiravir / BigDFT system (matrix dimension $147\times 147$). We examine the error for four different floating point types: BFLOAT16 (8, 7), FP16 (5, 10), AMD FP24 (7, 16), and FP16 with FP32 (8, 23) Accumulation (TC). The gray box extends from the NWChem convergence cutoffs down to two orders of magnitude lower. “Mixed Abs. Energy” labels the energy error after applying one McWeeny step in full double precision.
@@
In all cases, the scaling scheme seamlessly handles the exponent, and hence the precision comes down to the size of the mantissa. Using plain FP16, 14 to 16 splits (105 – 136 multiplications) is required. This is a substantial increase in the number of multiplications compared to the Tensor Core version, however future hardware may not offer FP16 TC (especially non-NVIDIA chips), with different performance ratios. For a fully converged double precision result on random matrices, 20 splits / 210 multiplications was required, leading again to an approximate factor of two savings. We compare as well the intermediate sized mantissa of AMD’s FP24, and find that it converges around 7 splits / 28 multiplications. For this system, the TRS4 algorithm required 26 calls to the Ozaki scheme multiplication routine to converge. Hence, when using FP16 TCs an overall speedup would require a time to solution ratio of 156 – 260 multiplications to one double precision diagonalization. For pure FP16, the requirement increases to 2756 – 3536 multiplications.
4.3 Potential to Exploit Sparsity
Depending on the underlying distribution of matrix values, the splitting procedure may introduce a significant amount of sparsity. Ichimura et al. 55 proposed to exploit the Ozaki scheme induced sparsity on a CPU by converting matrices with a sparsity level above 90% to the ELL format, and performing sparse matrix - dense matrix multiplication. Sparsity can be more directly exploited to accelerate calculations on Tensor Cores: on an A100 GPU, the FP16 TC performance increases from 312 TFLOPS to 624 TFLOPS for sparse matrices. This sparsity feature requires a special sparsity structure — for any given row / column of four elements, at least two of them must be zero.
@@ 4.3 Potential to Exploit Sparsity
We examined the final multiplication performed for the Molnupiravir system by the purification algorithm when using the Tensor Core representation (FP16 with FP32 accumulation) and four splits to see how frequently this sparsity pattern appeared. An entry of the split matrix (after scaling) was considered to be zero if it fell below the smallest value representable by double precision. The first split of both matrices have a number of non-zero blocks below 31%, however the next split is above 95%, and subsequent splits are nearly fully dense. For the pure FP16 implementation with 14 splits, the matrices are substantially more sparse: the number of non-zeros blocks is below 1% for the first split, 3% for the second split, and 9% for the third split. By the seventh split the sparsity of both matrices was above 59% (and for all subsequent splits). Thus we anticipate that some sparsity may be exploited in future work, though it would be limited to the first set of splits.
4.4 Basis Set and Size Effects
@@
Figure 8: Impact of the basis set on the number of splits required in the Ozaki scheme for the selenite in water system. Matrix sizes are: $179\times 179$ (PCSEG-0), $309\times 309$ (PCSEG-1), $515\times 515$ (AUG-PCSEG-1), $713\times 713$ (PCSEG-2), $1117\times 1117$ (AUG-PCSEG-2), and $1508\times 1508$ (PCSEG-3). All calculations are performed with FP16 Tensor Cores. The gray box extends from the NWChem convergence cutoffs down to two orders of magnitude lower. “Mixed Abs. Energy” labels the energy error after applying one McWeeny step in full double precision.
In Fig. 8 we plot the error for each of the basis sets with a given number of splits. We observe that for the larger basis sets, the required number of splits increases from four to five — particularly for the errors in the commutator and energy. This is consistent with our previous results where the difference between four and five splits was small, but not negligible. Overall, our earlier findings transfer well to even challenging numerical conditions. Thus, purification with the Ozaki scheme should be applicable to routinely performed quantum chemistry calculations, and not just specialized approximate schemes.
We also investigate the number of splits required as a function of system size using water clusters of increasing size computed with BigDFT. This analysis will allow us to separate the effect of basis set conditioning with matrix size. For the water clusters, the largest matrix is of dimension $1806\times 1806$, and for the selenite systems it was $1508\times 1508$ (PCSEG-3). The errors with respect to system size are plotted in Fig. 9. We find a modest growth in the number of splits required with system size. The error with five splits is similar for the largest water cluster as for the selenite system with the largest basis set, with the commutator and energy errors being about an order of magnitude lower. Thus, we conclude that five splits is a sufficient recommendation, though automated schemes may improve the usability of the Ozaki scheme.
Figure 9: Impact of the system size on the number of splits required in the Ozaki scheme for the water cluster systems. BigDFT uses 6 basis functions per water molecule, so that the smallest matrix is of dimension $30\times 30$ and the largest $1806\times 1806$. All calculations are performed with FP16 Tensor Cores. The gray box extends from the NWChem convergence cutoffs down to two orders of magnitude lower. “Mixed Abs. Energy” labels the energy error after applying one McWeeny step in full double precision.
4.5 Comparison with the Markidis Method
In the previous work of Finkelstein et al. that performed density matrix purification on NVIDIA Tensor Core units 45, 46, 47, the precision of the result was improved using the method of Markidis 2. Here, the matrix $X$ is split into a lower and high precision part:
@@ 4.5 Comparison with the Markidis Method
In the Markidis method, we can include up to four terms, each refining the precision of the result. In Tab. 2 we show the precision and the number of purification iterations required (which may increase due to low precision) for a given number of terms for various systems. We caution that the number of iterations may be made more uniform by a more sophisticated convergence test 73.
@@
RMSD Errors
Molnupiravir $2.2\times 10^{-4}$ $1.4\times 10^{-4}$ $5.1\times 10^{-7}$ $4.4\times 10^{-7}$ $2.1\times 10^{-7}$ $8.5\times 10^{-13}$ $7.0\times 10^{-15}$
Water 237 $6.1\times 10^{-5}$ $3.7\times 10^{-5}$ $2.0\times 10^{-7}$ $1.8\times 10^{-7}$ $1.1\times 10^{-7}$ $4.2\times 10^{-10}$ $8.1\times 10^{-15}$
Selenite/A $5.6\times 10^{-4}$ $3.2\times 10^{-4}$ $1.1\times 10^{-6}$ $1.1\times 10^{-6}$ $5.8\times 10^{-7}$ $6.5\times 10^{-11}$ $1.1\times 10^{-15}$
Selenite/B $3.8\times 10^{-4}$ $2.3\times 10^{-4}$ $1.2\times 10^{-6}$ $1.3\times 10^{-6}$ $8.6\times 10^{-7}$ $7.8\times 10^{-11}$ $1.3\times 10^{-15}$
Selenite/C $2.7\times 10^{-4}$ $1.6\times 10^{-4}$ $2.8\times 10^{-6}$ $2.9\times 10^{-6}$ $8.0\times 10^{-7}$ $1.4\times 10^{-9}$ $1.0\times 10^{-15}$
Commutator Errors
Molnupiravir $2.2\times 10^{-2}$ $1.3\times 10^{-2}$ $5.5\times 10^{-5}$ $4.6\times 10^{-5}$ $2.3\times 10^{-5}$ $8.6\times 10^{-11}$ $2.8\times 10^{-14}$
Water 237 $6.5\times 10^{-2}$ $4.1\times 10^{-2}$ $2.2\times 10^{-4}$ $1.9\times 10^{-4}$ $1.1\times 10^{-4}$ $4.7\times 10^{-7}$ $2.9\times 10^{-13}$
Selenite/A $4.0\times 10^{-1}$ $1.5\times 10^{-1}$ $4.2\times 10^{-3}$ $4.5\times 10^{-3}$ $4.3\times 10^{-3}$ $5.1\times 10^{-7}$ $4.1\times 10^{-13}$
Selenite/B $4.1\times 10^{-1}$ $1.6\times 10^{-1}$ $5.2\times 10^{-3}$ $5.7\times 10^{-3}$ $4.6\times 10^{-3}$ $9.5\times 10^{-7}$ $1.4\times 10^{-12}$
Selenite/C $1.3\times 10^{0}$ $6.6\times 10^{-1}$ $1.4\times 10^{-2}$ $1.6\times 10^{-2}$ $1.2\times 10^{-2}$ $4.3\times 10^{-5}$ $3.8\times 10^{-12}$
Abs. Energy Errors
Molnupiravir $2.6\times 10^{-3}$ $1.9\times 10^{-3}$ $8.7\times 10^{-5}$ $8.5\times 10^{-5}$ $1.2\times 10^{-5}$ $4.8\times 10^{-13}$ $1.1\times 10^{-13}$
Water 237 $1.9\times 10^{-2}$ $2.2\times 10^{-5}$ $1.2\times 10^{-3}$ $1.2\times 10^{-3}$ $3.6\times 10^{-4}$ $3.5\times 10^{-9}$ $1.7\times 10^{-12}$
Selenite/A $1.3\times 10^{-1}$ $1.4\times 10^{-3}$ $4.3\times 10^{-3}$ $4.3\times 10^{-3}$ $1.0\times 10^{-3}$ $5.4\times 10^{-10}$ 00
Selenite/B $5.8\times 10^{-2}$ $8.7\times 10^{-2}$ $8.5\times 10^{-3}$ $8.4\times 10^{-3}$ $2.2\times 10^{-3}$ $6.5\times 10^{-10}$ $4.5\times 10^{-13}$
Selenite/C $3.5\times 10^{-1}$ $2.1\times 10^{-1}$ $1.3\times 10^{-2}$ $1.3\times 10^{-2}$ $4.0\times 10^{-3}$ $6.3\times 10^{-8}$ $9.1\times 10^{-13}$
Mixed Abs. Energy Errors
Molnupiravir $1.9\times 10^{-4}$ $6.2\times 10^{-5}$ $1.2\times 10^{-9}$ $9.8\times 10^{-10}$ $2.0\times 10^{-10}$ $1.4\times 10^{-13}$ $1.1\times 10^{-13}$
Water 237 $1.6\times 10^{-3}$ $5.3\times 10^{-4}$ $2.3\times 10^{-8}$ $1.9\times 10^{-8}$ $6.9\times 10^{-9}$ $8.0\times 10^{-13}$ $1.7\times 10^{-12}$
Selenite/A $8.7\times 10^{-3}$ $2.4\times 10^{-3}$ $8.4\times 10^{-8}$ $8.3\times 10^{-8}$ $2.9\times 10^{-8}$ $1.6\times 10^{-11}$ 00
Selenite/B $8.8\times 10^{-3}$ $3.0\times 10^{-3}$ $2.4\times 10^{-7}$ $2.8\times 10^{-7}$ $6.9\times 10^{-8}$ $2.3\times 10^{-11}$ $4.5\times 10^{-13}$
Selenite/C $5.4\times 10^{-2}$ $1.6\times 10^{-2}$ $4.2\times 10^{-6}$ $4.6\times 10^{-6}$ $5.3\times 10^{-7}$ $7.5\times 10^{-11}$ $9.1\times 10^{-13}$
@@
Molnupiravir $17$ $27$ $27$ $24$ $27$ $13$ $13$
Water 237 $20$ $45$ $17$ $15$ $13$ $11$ $11$
Selenite/A $101$ $74$ $22$ $29$ $23$ $22$ $22$
Selenite/B $72$ $101$ $37$ $33$ $30$ $27$ $27$
Selenite/C $101$ $66$ $31$ $41$ $32$ $32$ $28$
@@
Table 2: Errors and number of purification iterations for the various systems computed with FP16 Tensor Cores vs. the numbers of correction terms. Non-converged systems are labeled as NC (not converged) in the number of iterations. A: PCSEG-1; B: AUG-PCSEG-1; C: PCSEG-3. 1 Term: $A^{0}B^{0}$; 2 Terms: $A^{0}B^{0}+A^{0}B^{1}$; 3 Terms: $A^{0}B^{0}+A^{0}B^{1}+A^{1}B^{0}$; 4 Terms: $A^{0}B^{0}+A^{0}B^{1}+A^{1}B^{0}+A^{0}B^{0}$. 4-MPFR: the 4 Term MPFR implementation with a double precision exponent. Ozaki (5): Ozaki scheme with 5 splits. FP64: calculations performed with double precision.
@@
While comparing our CUDA and MPFR implementations of the Markidis method, we noted that a significant amount of error was introduced due to the limited exponent range. To improve this, we applied the same scaling method for the exponent ranges used in the Ozaki scheme 56, and found it provided substantial improvement. For example, the 4 Terms result for Selenite/PCSEG-3 improved from an RMSD error of $1.72\times 10^{-5}$ to $2.86\times 10^{-6}$. Nonetheless, the MPFR result with a double precision exponent still remains more precise; in the future, it may be possible to further improve the Markidis method with a new scaling scheme.
@@
In the implementation of Finkelstein et al., they include three terms in the Markidis correction, which is well justified from our data. This three term result can be accomplished at the cost of only two multiplications because they implement a scheme that only requires squaring symmetric matrices, which means that one can exploit the relation $A^{1}A^{0}=(A^{0}A^{1})^{T}$ even in inexact arithmetic. The benefit of the Markidis method is thus that it can substantially improve the accuracy at a low cost with FP16 Tensor Cores (especially if combined with the exponent scaling scheme). On the other hand, it does not converge to the double precision result, can’t take advantage of lower precision hardware like FP16 without FP32 accumulation, and has significant errors in more challenging numerical conditions (like larger systems and basis sets). In particular, the commutator error remains large for the Markidis method, and for the Selenite/C system the energy error is above $10^{-6}$ Hartree even with the iterative refinement approach. For this reason, the higher order approximation of the Ozaki scheme is valuable. When only the matrix square is required, the Ozaki scheme can also exploit this symmetry in the calculation of the error free terms (see Fig. 2) as well as other optimizations detailed by Uchino 74. Using this symmetry and only computing the square of the matrix to purify, the Ozaki scheme with four or five terms would require 8 or 11 multiplications respectively (compared to the two required by the Markidis scheme).
4.6 Mixed Precision Purification